2. [0pt] In scuba diving, a regulator is used so that the pressure of the air the diver breathes is close to that of the ambient water. A reckless swimmer decides to use a hose sticking out of the surface to breathe underwater while diving in a lake. When the air pressure in the lungs is at a pressure of around 0.160 atm below the ambient pressure, lung injury may occur. Find the depth at which the swimmer would experience such a pressure differential.
Answer #1

The pressure as he dives increases as ρ*g*h

so for a pressure difference of 0.160atm = 0.160atm*1.01x10^5Pa/atm = 16160 Pa
the corresponding depth h is

h = 16160/(ρ*g) = 16160/(1000*9.8) = 1.648m